OpenStudy (anonymous):

Solve for x: |3x + 2| = 2 x = 0 and x = negative 4 over 3 x = 4 over 3 and x = negative 4 over 3 x = 0 No Solutions

OpenStudy (lgbasallote):

hint: solve for x in 3x + 2 = 2 and 3x + 2 = -2 does that help?

OpenStudy (anonymous):

3x + 2 = 2 3x = 0 x = 0 AND -(3x +2) = 2 -3x - 2 = 2 -3x = 4 x = -3/4

OpenStudy (anonymous):

the second option?

OpenStudy (cwrw238):

x = -4/3